# Bridging AI and Science: Implications from a Large-Scale Literature Analysis of AI4Science

### v1

@@ 6. Limitations, Challenges, and Future Work
Our analysis focuses on publications from leading science and AI venues, specifically top-tier journals and conferences. While this ensures high-quality and influential work, it introduces a selection bias that may overlook significant contributions from less prominent venues, potentially resulting in an incomplete representation of the AI4Science landscape. To mitigate this bias, future work could expand the dataset to include a broader range of publications. Given that our analysis framework is flexible and can easily accommodate new data sources, such an expansion would be a valuable direction for achieving a more comprehensive understanding of the AI4Science field.
Abstract-based extraction.
The extraction of scientific problems, AI methods, and AI usages in this study rely solely on paper titles and abstracts, due to limited access to full texts and cost considerations. This approach aligns with the existing AI4Science literature analysis efforts, which also focus on the title and abstract as in the metadata (Gao and Wang, 2023; Duede et al., 2024). While this method provides useful insights, it may miss important details present in the full text of publications. For example, when extracting the AI methods used in the AlphaFold paper (Jumper et al., 2021), as illustrated in Fig. 1, GPT identifies only the general use of a Neural Networks approach from the abstract. However, the specific use of the backbone Transformer model is detailed in the full text. In future work, efforts can be directed toward incorporating full-text analysis to enrich the extracted content and capture more nuanced AI methodologies used in scientific research.
The definition of “AI4Science”.
There is no universally accepted standard for defining the boundaries of the broad terms “AI”, “science”, and “AI4Science”, which presents a challenge for accurate classification. Previous studies have relied on keywords and human annotations (Gao and Wang, 2023; Duede et al., 2024), but this method can overlook less-known techniques and lacks the ability to capture the full semantic space of AI methods used in science. In contrast, our approach leverages the knowledge and capabilities of LLMs to classify AI4Science work, with human verification on a small-scale subset. To foster more consistent and scalable analyses, the AI4Science community should work toward developing standardized classification guidelines for automatic large-scale literature analysis.
@@
Our primary goal in this study is not to optimize the link prediction model, but rather to use link prediction as a lens to explore the potential of connecting AI methods and scientific problems. As a result, we only included a few link prediction models, leaving room for further research to improve the performance of AI4Science link prediction or building recommender systems for AI methods or scientific problems.
@@ 7. Conclusion
This paper aims to identify and bridge the gaps between AI and science in the context of AI4Science research. We introduce a comprehensive, large-scale dataset of AI4Science publications, with scientific problems and AI methods extracted using large language models (LLMs). Through quantitative analysis of this dataset, we uncover several key disparities: (1) Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods under-investigated; (2) The connectivity between scientific problems and AI methods is highly imbalanced, with certain nodes acting as “hubs” while other peripheral nodes are less connected; (3) The science and AI communities take distinct approaches to integrating AI into scientific research, prioritizing different problems and methods. We further investigate the potential and challenges of fostering AI4Science collaboration through the lens of link prediction. The experiment results demonstrate great opportunities for using link prediction models to explore the under-investigated scientific problems and AI methods, as well as discover novel connections. We anticipate that our findings and tools will provide valuable insights to enhance interdisciplinary collaboration and accelerate scientific discovery through the integration of AI.
Acknowledgements.
The authors would like to extend their sincere appreciation to Shixuan Liu and Sicheng Shen for their contributions as annotators.
@@ A.1. Curation of Literature
  • For AI domain, we obtained our data from MASSW (Zhang et al., 2024), which encompasses the following six top AI conferences: AAAI, IJCAI, ICLR, ICML, NeurIPS, WWW. Only papers with both title and abstract available and within 2014 to 2024 are kept. A brief statistics of AI domain papers are as follows:
@@ A.3. Semantic Clustering
After obtaining the six key aspects of the papers, we first leverage the InstructionEmbedding model(instructor large (Su et al., 2022)) to obtain the semantic embeddings of the scientific problems and AI methods extracted.
@@
After we obtained the semantic embeddings, we used LargeVis (Tang et al., 2016) for dimension reduction followed by HDBSCAN (Campello et al., 2013) for clustering.
Summarization
@@
You are an assistant in scientific research. Given a cluster of scientific problems/challenges, please help summarizing the cluster into a keyword or keyphrase less than 3 words. The top appearing words in the cluster together with some examples of extracted paper from that cluster will be provided. Please summarize based on the provided information.
@@
## Examples of scientific papers
Below are some specific examples of scientific challenges in this cluster. For each sample the problem keyword, definition and discipline will be provided.
@@
Please output the summarization as a list, which has exactly *one* elements. The output should be in the format of a python list.
["Keyword"]
@@
You are an assistant in Artificial Intelligence research, where where AI generally refers to intelligence exhibited by machines (particularly computer systems), including models, algorithms. Given a cluster of Artificial Intelligence methods, please help summarizing the cluster into a keyword or keyphrase less than 3 words. The top appearing words in the cluster together with some examples of Artificial Intelligence metho from that cluster will be provided. Please summarize based on the provided information.
@@
Below are some specific examples of Artificial Intelligence methods in this cluster. For each sample the problem keyword, definition and discipline will be provided.
@@
Please output the summarization as a list, which has exactly *one* elements. The output should be in the format of a python list.
["Keyword"]
Merging Clusters
To refine previous clustering results and obtain more consistent clusters, We perform cluster merging based on LLM summarization. If LLM gives the same summary for a group of adjacent clusters on the 2D projection map, we merge them into a large cluster with the same summary. After merging, we have $369$ scientific problems clusters and $345$ AI methods clusters as the final clusters for further analysis.
@@ Appendix B Discrepancy Analysis
Bipartite graph node degrees.
Table 12 presents the nodes with the highest degrees in the bipartite graph, representing the most connected scientific challenges and AI methods in the AI4Science landscape. These high-degree nodes act as key “hubs,” indicating their central role in bridging AI techniques with scientific problems
Discrepancies between the AI and science communities.
Fig. 4-5 illustrates the distribution of AI4Science and non-AI4Science publications across different communities, highlighting how AI and scientific fields engage with interdisciplinary research distinctively. Table.
@@
1 Neural Network Challenges 229 Machine Learning 218
2 Statistical Inference 162 Deep Learning 149
3 Language and Auditory Processing 149 Genomic Analysis 122
4 Optimization Problems 135 Advanced Imaging and Spectroscopy 91
5 Social Media Dynamics 134 Computational Biology 79
6 Decision-Making Processes 126 Protein Design 76
7 Market Economics 118 Neural Networks 70
8 Game Theory 115 Molecular Modeling 65
9 Dynamical Systems 115 Image Processing 64
10 Imaging Challenges 106 Classification Methods 60
11 Causal Inference 105 Data Analysis 58
12 Visual and Sensory Perception 97 Bayesian Inference 57
13 Robotic Actuators 92 Materials Informatics 55
14 Urban and Sustainability Challenges 84 Unsupervised Learning 52
15 Optical and Wireless Communication 79 Wearable Sensor Technology 51
16 Complex Network Dynamics 78 Predictive Modeling 51
17 Scientific Policy Challenges 76 Drug Screening 49
18 Neural Imaging 72 Deep Neural Networks 49
19 Traffic Prediction 69 Convolutional Neural Networks 46
20 Protein Structure and Design 63 Artificial Intelligence Techniques 44
@@
1 Neural Network Challenges Neural Network Challenges Machine Learning Deep Learning
2 Imaging Challenges Social Media Dynamics Deep Learning Machine Learning
3 Language and Auditory Processing Statistical Inference Genomic Analysis Data Analysis
4 Visual and Sensory Perception Language and Auditory Processing Advanced Imaging and Spectroscopy Neural Networks
5 Protein Structure and Design Game Theory Protein Design Game Theory
6 Materials Design Market Economics Neural Networks Bayesian Inference
7 Artificial Intelligence Ethics Optimization Problems Computational Biology Causal Inference
8 Statistical Inference Decision-Making Processes Molecular Modeling Probabilistic Modeling
9 Single-Cell RNA Sequencing Dynamical Systems Optical Computing Reinforcement Learning
10 Single-Cell Omics Causal Inference Materials Informatics Optimization Methods
11 Decision-Making Processes Robotic Actuators Image Processing Predictive Modeling
12 Cancer Detection Imaging Challenges Neuroimaging Techniques Sequential Decision Making
13 Gene Expression Regulation Optical and Wireless Communication Memory-augmented Systems Convolutional Neural Networks
14 Genetic Variants Traffic Prediction Drug Screening Spatio-temporal Analysis
15 Dynamical Systems Urban and Sustainability Challenges Wearable Sensor Technology Fair Resource Allocation
16 Chemical Reaction Dynamics Fair Resource Allocation Unsupervised Learning Regret Minimization
17 Protein-Protein Interactions Voting and Elections Molecular Dynamics Policy Optimization
18 Wearable Tactile Sensors Visual and Sensory Perception Classification Methods Auction Mechanism Design
19 Antibiotic Resistance Scientific Policy Challenges Deep Neural Networks Classification Methods
20 Motor Control and Learning Complex Network Dynamics Bayesian Inference Generative Models
@@ C.1. Experiment Setups
Number of All publications 138,577 20,718
AI4Science publications 5,829 1,195
Pub. in well-investigated sci. clusters 3,503 512
Pub. in under-investigated sci. clusters 1,129 379
Pub. in well-investigated AI clusters 3,038 687
Pub. in under-investigated AI clusters 1,336 186
Avg. deg. of Scientific problem nodes 14.4 3.7
Well-investigated sci. nodes 59.5 8.9
Under-investigated sci. nodes 5.3 2.5
AI method nodes 14.7 4.7
Well-investigated AI nodes 25.6 7.4
Under-investigated AI nodes 7.8 1.8
@@
  • Bipartite graph link prediction (“$P_{i}\to M_{i}$” or “$M_{i}\to P_{i}$”): In this approach, scientific problems and AI methods are represented as high-level cluster labels, serving as nodes in a bipartite graph (as defined in Sec. LABEL:). This captures more generalized connections between similar scientific problems and AI methods.
  • Generative link prediction (“$p_{i}\to m_{i},u_{i}$” or “$m_{i}\to p_{i},u_{i}$”): Here, scientific problems and AI methods are expressed in detailed text descriptions rather than coarse cluster labels. This approach leverages textual generative models, such as large language models (LLMs), to capture more fine-grained distinctions and predict links based on textual data.
@@
For bipartite link prediction, we evaluate performance using Recall, Precision, and F1 scores @K, where K represents the number of generated candidate links, as is standard in the literature (Al Hasan et al., 2006; Yang et al., 2015). Particularly, for a source node $s$, supposing in the test data, it is connected to a set of target nodes $T(s)$, and the model generates $K$ predictions $P(s)$. The metrics are calculated as below:
@@
For link generation, we adopt three widely used evaluation metrics for text generation: ROUGE (Lin, 2004), BLEURT (Sellam et al., 2020), and the cosine similarity of text embeddings. The best scores @K are reported for each of these metrics.
@@
LLM generation setup
For generative link prediction, we consider both Sci $\to$ AI and AI $\to$ Sci generations:
@@
For generative link prediction with LLMs (LLM+RAG), we use the following prompt [^1] For brevity only Sci $\to$ AI RAG prompt is provided:
@@
For graph-level link prediction with LLMs (LLM+Graph), we use the following prompt:  [^2] For brevity only Sci $\to$ AI RAG prompt is provided:

### v2

@@ 6. Limitations, Challenges, and Future Work
As an emerging and interdisciplinary field, cutting-edge and influential AI4Science studies often originate from prominent venues of general interest. For instance, groundbreaking work such as the AlphaFold (Jumper et al., 2021) was published in Nature. Based on this observation, we have focused on top-tier venues from the last decade. While this ensures high-quality and impactful contributions, it could introduce a selection bias that may overlook significant contributions from less prominent venues, potentially resulting in an incomplete representation of the AI4Science landscape. Given the flexibility of our analysis framework, researchers can integrate additional data sources from less-prominent venues to achieve a more representative view.
LLM-based extraction from titles and abstracts.
The LLM-based extraction in this study relies solely on paper titles and abstracts, due to limited access to full texts, challenge of processing PDF formats, relatively low signal-to-noise ratio in full texts, and cost considerations. Our approach is comparable to the existing AI4Science literature analysis efforts, which also focus on the title and abstract as in the metadata (Gao and Wang, 2023; Duede et al., 2024). While key research problems and methods are often described in abstracts, and human evaluation confirms the accuracy of LLM extraction (Appendix A.2), we may miss details presented in the full text of publications. Future work can build upon our general and flexible framework to incorporate full-text analysis, enriching the extracted content and capturing more nuanced AI methodologies used in scientific research.
@@
Our goal is not to optimize the link prediction model, but rather to use link prediction as a lens to demonstrate the potential of connecting AI methods and scientific problems. As a result, we only included a few straightforward link prediction models, leaving room for further research to improve the performance of AI4Science link prediction or building recommender systems for AI methods or scientific problems.
@@ 7. Conclusion
This paper aims to depict the AI4Science landscape, identify current gaps between AI and science, while exploring potential ways to bridge them. We introduce a comprehensive, large-scale dataset of curated AI4Science publications, with scientific problems and AI methods extracted using large language models. Through quantitative analysis of this dataset, we uncover several key disparities: (1) Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods underexplored; (2) The connectivity between scientific problems and AI methods is highly imbalanced, with certain nodes acting as “hubs” while other peripheral nodes are less connected; (3) The science and AI communities take distinct approaches to integrating AI into scientific research, prioritizing different problems and methods. We further investigate the potential and challenges of fostering AI4Science collaboration through the lens of link prediction. The experiment results demonstrate great opportunities for using link prediction models to explore the under-investigated scientific problems and AI methods, as well as discover novel connections. We anticipate that our findings and tools will provide valuable insights to enhance interdisciplinary collaboration and accelerate scientific discovery through the integration of AI.
@@ A.1. Curation of Literature
  • For AI domain, we obtained our data from MASSW (Zhang et al., 2024), which encompasses the following seven top AI conferences: AAAI, IJCAI, ICLR, ICML, NeurIPS, WWW, SIGKDD. Only papers with both title and abstract available and within 2014 to 2024 are kept. A brief statistics of AI domain papers are as follows:
@@ results
The high extraction accuracy indicated by our human evaluation also aligns with human evaluation results from recent studies (Zhang et al., 2024), where LLMs have been used to extract scientific workflows from titles and abstracts.
@@ A.3. Semantic Clustering
After obtaining the six key aspects of the papers, we first leverage the InstructionEmbedding model (instructor large)(Su et al., 2022) to obtain the semantic embeddings of the scientific problems and AI methods extracted. InstructorEmbedding is an advanced embedding model that generates task- and domain-specific text embeddings without additional training. Given our need to embed both scientific problems and AI methods, we use InstructorEmbedding with instructions below to provide contextualized embeddings tailored to our task rather than context-free embedding methods
@@
After we obtained the semantic embeddings, we used LargeVis (Tang et al., 2016) for dimension reduction followed by HDBSCAN (Campello et al., 2013) for clustering. HDBSCAN is a widely used density-based clustering algorithm designed for practical data mining tasks. As an extension of DBSCAN, it employs hierarchical clustering to identify clusters of varying densities and is more robust to parameter selection. A density-based clustering method allows us to identify clusters at appropriate granularities and avoids a top-down decision to predefine the number of clusters. Its efficiency ($O(n\log n)$ complexity) also makes it suitable for our dataset size.
Labeling
@@
Using LLMs with the below prompts to summarize clusters encourage the cluster labels to have the desirable properties (Mei et al., 2007): (1) understandable, (2) semantically relevant, (3) well covering the whole topic, and (4) distinguishable from other topics.
@@
You are an assistant in scientific research. Given a cluster of scientific problems/challenges, please help summarizing the cluster into a keyword or keyphrase less than 3 words. Moreover, top words from several nearby clusters are provided and your summary should accurately capture the essence of the cluster while differentiating it from neighboring clusters. Please summarize based on the provided information.
@@
## Examples of Scientific Problems
Below are some specific examples of scientific problems in this cluster. For each sample the problem keyword, definition and discipline will be provided.
@@
## Top words from nearby clusters’ texual information
{ Nearby clusters’ top words }
@@
* The summarization should be specific enough to differentiate it from nearby clusters
* Also provide summarization for nearby clusters
@@
Please output the summarization as a list, which has exactly *(#nearby clusters+1) outputs. The output should be in the format of a python list.
["Keyword", "Keyword neighbor 1",…]
@@
You are an assistant in Artificial Intelligence research, where where AI generally refers to intelligence exhibited by machines (particularly computer systems), including models, algorithms. Given a cluster of Artificial Intelligence methods, please help summarizing the cluster into a keyword or keyphrase less than 3 words. The top appearing words in the cluster together with some examples of Artificial Intelligence methods from that cluster will be provided. Moreover, top words from several nearby clusters are provided and your summary should accurately capture the essence of the cluster while differentiating it from neighboring clusters.. Please summarize based on the provided information.
@@
Below are some specific examples of Artificial Intelligence methods in this cluster. For each sample the method keyword and definition will be provided.
@@
## Top words from nearby clusters’ texual information
{ Nearby clusters’ top words }
@@
* The summarization should be specific enough to differentiate it from nearby clusters
* Also provide summarization for nearby clusters
@@
Please output the summarization as a list, which has exactly *(#nearby clusters)* outputs. The output should be in the format of a python list.
["Keyword", "Keyword neighbor 1",…]
@@ Appendix B Discrepancy Analysis
Fig. 5-6 illustrates the distribution of AI4Science and non-AI4Science publications across different communities, highlighting how AI and scientific fields engage with interdisciplinary research distinctively.
@@
Well- and under-explored regions
Table 12 list some top clusters of well- and under-explored scientific problems as well as AI methods.
Scientific problems AI Methods
Well-explored Under-explored Well-explored Under-explored
Advanced Imaging Challenges Asymmetric Synthesis Deep Learning Models Attention Mechanisms
Visual Perception and Neuroscience Magnetic Materials and Spintronics Convolutional Neural Networks Gradient Descent Methods
Decision Making Forests and Climate Change Genomic Analysis Gradient-Based Methods
Neural Network Challenges Viral Mechanisms Neural Networks Policy Gradient Methods
Urban Traffic Management Evolutionary Paleontology Classification Methods Graph Embedding
Alzheimer’s Disease Mechanisms Nanostructure Synthesis Computational Modeling Regularization Methods
Statistical Inference Diabetes and Insulin Regulation Molecular Dynamics Simulations Transformer Architectures
Depression and Anxiety Bacterial Infections and Immune Response Predictive Modeling Contrastive Learning
Language Processing Cancer Signaling Pathways Advanced Imaging Techniques Graph Representation
Genetic Variants and Traits Cancer Drug Resistance Computational Drug Screening Adaptive Algorithms
Table 12. Scientific problems and AI methods that are well- and under-explored by AI4Science research. Cluster names are ranked by the number of total publications. Only top 10 clusters are shown.
Bipartite graph node degrees.
Table 13 presents the nodes with the highest degrees in the bipartite graph, representing the most connected scientific challenges and AI methods in the AI4Science landscape. These high-degree nodes act as key “hubs,” indicating their central role in bridging AI techniques with scientific problems.
Discrepancies between the AI and science communities.
@@
1 Neural Network Challenges 164 Machine Learning 140
2 Urban Traffic Management 122 Genomic Analysis Methods 137
3 Electronic Health Data Challenges 102 Advanced Imaging Techniques 98
4 Imaging Challenges 89 Computational Drug Screening 97
5 Statistical Inference 89 Molecular Dynamics Simulations 88
6 Decision Making 86 Protein Structure Design 88
7 Language Processing 85 Deep Learning Models 84
8 Dynamical Systems 84 Neural Networks 62
9 Visual Perception and Neuroscience 82 Classification Methods 62
10 Market Pricing and Allocation 77 Molecular Structural Design 61
11 Neural Imaging Challenges 74 Deep Neural Networks 57
12 Single-cell RNA Sequencing 71 Predictive Modeling 53
13 Soft Robotics 69 Convolutional Neural Networks 52
14 Drug Discovery 67 Computational Modeling 52
15 Protein Design and Function Prediction 67 Unsupervised Learning 50
16 Social Media Dynamics 67 Image Segmentation 50
17 Mechanical Properties of Materials 64 Data Analysis 41
18 Causal Inference 62 Deep Learning Framework 39
19 Optimization Problems 60 Data-driven Approaches 39
20 Game Theory Challenges 58 Recurrent Neural Networks 37
@@
1 Neural Network Challenges Urban Traffic Management Machine Learning Deep Learning Models
2 Imaging Challenges Neural Network Challenges Genomic Analysis Methods Machine Learning
3 Single-cell RNA Sequencing Electronic Health Data Challenges Deep Learning Models Reinforcement Learning
4 Mechanical Properties of Materials Language Processing Molecular Dynamics Simulations Attention Mechanisms
5 Visual Perception and Neuroscience Statistical Inference Advanced Imaging Techniques Convolutional Neural Networks
6 Protein Design and Function Prediction Market Pricing and Allocation Computational Drug Screening Neural Networks
7 Cancer Detection Soft Robotics Protein Structure Design Recurrent Neural Networks
8 Genetic Variants and Traits Dynamical Systems Molecular Structural Design Graph Neural Networks
9 Decision Making Decision Making Brain Imaging Techniques Policy Gradient Methods
10 Language Processing Game Theory Challenges Computational Modeling Multitask Learning
11 Gene Regulatory Networks Causal Inference Optical AI Methods Classification Methods
12 Drug Discovery Drug Discovery Deep Neural Networks Transformer Architectures
13 Cancer Mutation Studies Optimization Problems Wireless Sensor Technology Spatio-Temporal Analysis
14 Mitochondrial DNA Mutations Social Media Dynamics Neural Networks Deep Reinforcement Learning
15 Protein-Protein Interactions Imaging Challenges Image Segmentation Predictive Modeling
16 Neural Imaging Challenges Urban Dynamics Unsupervised Learning Causal Inference
17 Cryo-electron microscopy Neural Imaging Challenges Classification Methods Self-Supervised Learning
18 Tactile Sensing Technology Visual Perception and Neuroscience Convolutional Neural Networks Graph Convolutional Networks
19 Chemical Reaction Dynamics Sparse Signal Processing Predictive Modeling Unsupervised Learning
20 Chromatin Organization Energy Management Neuromorphic Devices Latent Variable Models
@@ C.1. Experiment Setups
Number of All publications 141,639 21,017
AI4Science publications 6,287 1,225
Pub. in well-explored sci. clusters 4,262 594
Pub. in under-explored sci. clusters 1,180 392
Pub. in well-explored AI clusters 3,076 742
Pub. in under-explored AI clusters 1,387 198
Avg. deg. of Scientific problem nodes 11.3 2.8
Well-explored sci. nodes 37.7 6.9
Under-explored sci. nodes 3.9 1.4
AI method nodes 12.4 3.0
Well-explored AI nodes 16.2 4.9
Under-explored AI nodes 7.2 1.3
@@
  • Cluster-level link prediction (“$P_{i}\to M_{i}$” or “$M_{i}\to P_{i}$”): In this approach, scientific problems and AI methods are represented as high-level cluster labels, serving as nodes in a bipartite graph. This captures more generalized connections between similar scientific problems and AI methods.
  • Paper-level link prediction (“$p_{i}\to m_{i},u_{i}$” or “$m_{i}\to p_{i},u_{i}$”): Here, scientific problems and AI methods are expressed in detailed text descriptions rather than coarse cluster labels. This approach leverages textual generative models, such as large language models (LLMs), to capture more fine-grained distinctions and predict links based on textual data.
@@
For cluster-level link prediction, we evaluate performance using Recall, Precision, and F1 scores @K, where K represents the number of generated candidate links, as is standard in the literature (Al Hasan et al., 2006; Yang et al., 2015). Particularly, for a source node $s$, supposing in the test data, it is connected to a set of target nodes $T(s)$, and the model generates $K$ predictions $P(s)$. The metrics are calculated as below:
@@
For paper-level link prediction, we adopt three widely used evaluation metrics for text generation: ROUGE (Lin, 2004), BLEURT (Sellam et al., 2020), and the cosine similarity of text embeddings. The best scores @K are reported for each of these metrics.
@@
LLM (Paper) link prediction setup
For paper-level generative link prediction, we consider both Sci $\to$ AI and AI $\to$ Sci generations:
@@
For paper-level link prediction with LLMs, we use the following prompt [^1] For brevity only Sci $\to$ AI RAG prompt is provided:
@@
For cluster-level link prediction with LLMs (LLM (Cluster)), we use the following prompt:  [^2] For brevity only Sci $\to$ AI RAG prompt is provided: